# Accelerated Gradient Methods with Biased Gradient Estimates: Risk Sensitivity, High-Probability Guarantees, and Large Deviation Bounds

### v1

@@ 5. Main Results for Strongly Convex Objectives
Recent work by [57] establishes a $4\times 4$ MI which, if satisfied, guarantees both a finite upper bound on the $H_{\infty}$-norm and a descent property for the Lyapunov function $V_{P,c_{1}}(\cdot)$. This result is tight in the sense that, for certain parameter choices, the bound on the $H_{\infty}$-norm matches the corresponding quadratic bounds up to some constants, and in the noiseless case, this MI recovers the rate results implied by the $3\times 3$ MI provided in [12]. We include this result in the appendix (theorem˜A.1) for completeness, together with its corollaries that provide explicit bounds $\overline{H}_{\infty}$ on the $H_{\infty}$ norm corresponding to GD, NAG-beta-opt and NAG-pop which are obtained by constructing explicit solutions to the MI. In the next result, we show that when the $4\times 4$ MI holds, it also enables us to derive bounds on the risk sensitivity $R_{K}(\theta)$ for any horizon $K$ and for any $\theta>0$ satisfying $\sqrt{\theta}~\overline{H}_{\infty}<1$. This type of an upper bound on $\theta$ to guarantee finiteness of $H_{\infty}$ is expected; a similar constraint arises in the quadratic case where computations are exact as in (4.12). The proof of this result mirrors the dynamic programming technique employed within the risk-sensitive control [62]. Since the risk metric $R_{K}(\theta)$ aggregates risk across iterations and the noise sequence $\{w_{k+1}\}_{k\geq 0}$ is conditionally independent from $\mathcal{F}_{k}$, we can isolate the contribution of each step via a backward–induction argument while fixing the iteration budget to $K$. This induction argument makes the critical parameter condition explicit, i.e., $\sqrt{\theta}\,\overline{H}_{\infty}<1$ that guarantees finiteness of the risk for all iteration budgets. The proof relies on a novel Lyapunov function, obtained by modifying $V_{P,c_{1}}(\xi_{k})$ into a weighted form that incorporates convex averages with $V_{P,c_{1}}(\xi_{k-1})$. Full details are provided in the appendix. This modification is crucial for deriving tighter bounds.
@@
Consider the noisy GMM iterations stated in (3.8) for minimizing $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$, starting from ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{0}=x_{-1}}\in\mathbb{R}^{d}$ with fixed parameters $(\alpha,\beta,\nu)$. Let ˜3.1 hold and consider the risk sensitive cost in (3.10) for some given arbitrary iteration budget $K\geq 1$ that is fixed.
@@
For GD with $\alpha\in(0,2/L)$, given any $K\geq 0$ and $\alpha\in(0,\frac{2}{L})$, the risk sensitive cost admits the following bound:
@@
$\overline{R}_{K}^{(2)}(\theta)$ $\coloneqq\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K+1}}\left[\frac{1-\left[1-2\mu\alpha(1-\frac{L\alpha}{2})+\alpha\mu|1-\alpha L|s_{\scriptscriptstyle\mathrm{GD}}\right]^{K+1}}{2\mu\alpha(1-\frac{L\alpha}{2})-\alpha\mu|1-\alpha L|{\scriptscriptstyle\mathrm{GD}}}\left[f(x_{0})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\right]\right$.
@@
For NAG with $\alpha\in(0,1/L]$ and $\beta=\nu=\frac{1-\sqrt{\alpha\mu}}{1+\sqrt{\alpha\mu}}$, the bound in theorem˜5.1 holds with constants
@@
$q=\frac{2}{s_{1}}\left(\frac{\alpha(1-\sqrt{\alpha\mu})}{4(1+\sqrt{\alpha\mu})^{2}}\right)+\frac{4L^{3}\alpha^{2}\nu^{2}}{2\mu s_{2}}\,,\quad r=a+\alpha^{2}\left(\frac{c_{1}L}{2}+\tilde{P}_{11}\right)=s_{1}+\frac{L\alpha^{2}(2+s_{2})}{2s_{2}}+\frac{\alpha}{2},$
@@
where we used the definition of $\bar{c}(K+1,\theta)$ and $a_{j,K+1}$ given in theorem˜5.1. Since the sequence $\{a_{1,j}\}_{j\geq 1}$ is non-decreasing in $j$ admitting $\lim_{j\to\infty}a_{1,j}=\frac{1}{1-(p+q)}$; the sequence $\ell_{j}:=\log\left(\frac{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})+r~\theta a_{1,j}}{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})-r~\theta a_{1,j}}\right)$ is also non-decreasing and admits the limit $\ell_{\infty}:=\lim_{j\to\infty}\ell_{j}=\log\left(\frac{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})+r~\theta 1/(1-(p+q))}{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})-r~\theta 1/(1-(p+q))}\right)=\log\left(\frac{1+\theta(\overline{H}_{\infty})^{2}}{1-\theta(\overline{H}_{\infty})^{2}}\right)$ using the definition of $\overline{H}_{\infty}$ from theorem˜5.1. Hence we conclude from (5.14) that
@@
Since $R(\theta)=\limsup_{K\to\infty}R_{K}(\theta)\leq\limsup_{K\to\infty}\overline{R}_{K}(\theta)$ and $R_{K}(\theta)=+\infty$ for $\sqrt{\theta}\,\overline{H}_{\infty}\geq 1$ for every $K$, we obtain (5.9) as desired. With similar computations, taking limit superior on both sides in (5.7) with respect to $K$ and using (5.8), we obtain $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq 4\sigma^{2}(\overline{H}_{\infty})^{2}.$ Finally, the fact that $\lim_{\theta\downarrow 0}\overline{R}(\theta)=4\sigma^{2}(\overline{H}_{\infty})^{2}$ follows simply by taking limit as $\theta\to 0$ in (5.9).
@@
For momentum methods with constant step size and unbiased stochastic gradient estimates, ergodic averages converge to a neighborhood controlled by the variance proxy $\sigma^{2}$. For example, for NAG-$\beta$-opt (using parameters as in table˜1) with i.i.d. noise $w_{k+1}\sim\mathcal{N}(0,\tfrac{\sigma^{2}}{d})$, [53, Cor. 4.8] gives $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\tfrac{\sqrt{\alpha}}{2\sqrt{\mu}}(1+\alpha L)$ for $\alpha\in(0,1/L]$, which vanishes as $\alpha\to 0$. In contrast, under biased noise the neighborhood is generally bounded away from zero due to a non-vanishing worst-case drift. From (5.10), the asymptotic error is governed by the bound on $H_{\infty}$ norm. For NAG-$\beta$-opt, under the ˜3.1, by corollary˜A.3 in the appendix, we have $\overline{H}_{\infty}=\tfrac{2\sqrt{5}}{\sqrt{\mu}}+\mathcal{O}(\sqrt{\alpha})$ as $\alpha\to 0$, and by (5.10) this yields $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\bigl{(}\tfrac{80}{\mu}+\mathcal{O}(\sqrt{\alpha})\bigr{)}$, and the right-hand side remains strictly positive as $\alpha\to 0$. Corollary 5.5 shows that momentum algorithms with smaller bound $\bar{H}_{\infty}$ exhibit tighter bounds on the asymptotic (suboptimality) bias, i.e., $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]$. Similar observations can be made for GD, HB or GMM. To our knowledge, this connection to the $H_{\infty}$-norm is entirely new for momentum methods.
For quadratic functions, we obtained large deviation results which obtains both the upper bound and the lower bound in (4.16) through an exact computation of the risk sensitive index as $K\to\infty$. Beyond quadratics, we have only bounds on the risk sensitive index. In the following result, we leverage this bound to obtain a large-deviation upper bound and high probability bounds for finite $K$.
@@
Under the premise of Theorem 5.1, consider the averaged iterates $\bar{x}_{K}=\frac{\sum_{j=0}^{K+1}x_{j}}{K+1}$. Then, for any $K\geq 1$ given and for any $\theta\in[0,~1/\overline{H}_{\infty}^{2})$ and $t\geq 0$, we have
@@
The first inequality in (5.15) is a consequence of the convexity of $f$. The second inequality is a Chernoff-type bound. More specifically, for $t\geq 0$, and $\theta\in(0,1/(\overline{H}_{\infty})^{2})$,
@@
where we used $\overline{R}_{K}(\theta)\geq R_{K}(\theta)$. Then, (5.16) is obtained by taking supremum over $\theta$. Finally, the lower bound on $\bar{I}_{K}(t)$ given in (5.17) is a consequence of Lemma E.4.
@@
$\bar{I}(t){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=}\begin{cases}\displaystyle\frac{t}{2\sigma^{2}(\overline{H}_{\infty})^{2}}\sqrt{1-\frac{4\sigma^{2}(\overline{H}_{\infty})^{2}}{t}}-\log\left(\frac{1+\sqrt{1-\frac{4\sigma^{2}(\overline{H}_{\infty})^{2}}{t}}}{1-\sqrt{1-\frac{4\sigma^{2}(\overline{H}_{\infty})^{2}}{t}}}\right),&\text{if}\quad t\geq 4\sigma^{2}(\overline{H}_{\infty})^{2};\\ 0,&\text{if}\quad t<4\sigma^{2}(\overline{H}_{\infty})^{2}.\end{cases}$ (5.20)
@@
By Corollary 5.5, $\limsup_{K\to\infty}\,\mathbb{E}\!\left[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\right]=\overline{m}:=4\sigma^{2}(\overline{H}_{\infty})^{2}.$ Since this value represents the asymptotic expected suboptimality, the events with suboptimality below $\overline{m}$ are not exponentially unlikely; hence, $\bar{I}(t)=0$ whenever $t<\overline{m}$, in accordance with (5.20). In contrast, for $t>\overline{m}$, (5.20) shows that $\bar{I}(t)$ is strictly decreasing in $\overline{H}_{\infty}$ over the region where $\bar{I}(t)>0$, i.e., when $0<\overline{H}_{\infty}<\tfrac{\sqrt{t}}{2\sigma}$. Therefore, smaller values of $\overline{H}_{\infty}$ correspond to strictly larger decay rates $\bar{I}(t)$ at the threshold $t$, provided that $t$ exceeds the critical level $\overline{m}=4\sigma^{2}(\overline{H}_{\infty})^{2}$. This shows that a smaller bound on $H_{\infty}$ norm—reflecting greater robustness to worst-case deterministic noise—leads to sharper concentration properties and faster decay of the tail probabilities associated with $f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}$. This is broadly consistent with the numerical behavior observed in LABEL:fig:rate-fn-nag-beta-opt-vary-stepsize for quadratics: reducing the stepsize increases the rate function $I(t)$ while simultaneously decreasing the bound on $H_{\infty}$-norm for NAG-beta-opt, implied by (A.8) in the appendix.

### v2

@@ 5. Main Results for Strongly Convex Objectives
Recent work by [57] establishes a $4\times 4$ MI which, if satisfied, guarantees both a finite upper bound on the $H_{\infty}$-norm and a descent property for the Lyapunov function $V_{P,c_{1}}(\cdot)$. This result is tight in the sense that, for certain parameter choices, the bound on the $H_{\infty}$-norm matches the corresponding quadratic bounds up to some constants, and in the noiseless case, this MI recovers the rate results implied by the $3\times 3$ MI provided in [12]. We include this result in the appendix (theorem˜A.1) for completeness, together with its corollaries (corollaries˜A.2 and A.3) that provide explicit bounds $\overline{H}_{\infty}$ on the $H_{\infty}$-norm corresponding to GD, NAG-beta-opt and NAG-pop which are obtained by constructing explicit solutions to the MI. In the next result, we show that when the $4\times 4$ MI holds, it also enables us to derive bounds on the risk sensitivity $R_{K}(\theta)$ for any horizon $K$ and for any $\theta>0$ satisfying $\sqrt{\theta}~\overline{H}_{\infty}<1$. This type of upper bound on $\theta$ to guarantee finiteness of $H_{\infty}$ is expected; a similar constraint arises in the quadratic case where computations are exact as in (4.12). The proof of this result mirrors the dynamic programming technique employed within the risk-sensitive control [62]. Since the risk metric $R_{K}(\theta)$ aggregates risk across iterations and the noise sequence $\{w_{k+1}\}_{k\geq 0}$ is conditionally independent from $\mathcal{F}_{k}$, we can isolate the contribution of each step via a backward–induction argument while fixing the iteration budget to $K$. This induction argument makes the critical parameter condition explicit, i.e., $\sqrt{\theta}\,\overline{H}_{\infty}<1$ that guarantees finiteness of the risk for all iteration budgets. The proof relies on a novel Lyapunov function, obtained by modifying $V_{P,c_{1}}(\xi_{k})$ into a weighted form that incorporates convex averages with $V_{P,c_{1}}(\xi_{k-1})$. Full details are provided in the appendix. This modification allows us to derive tighter bounds.
@@
Consider the noisy GMM iterations stated in (3.8) for minimizing $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$, starting from ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{0}=x_{-1}}\in\mathbb{R}^{d}$ with fixed parameters $(\alpha,\beta,\nu)$. Let ˜3.1 hold and consider the risk-sensitive cost in (3.10) for some given arbitrary iteration budget $K\geq 1$ that is fixed.
@@
For GD, given any $K\geq 0$ and $\alpha\in(0,\frac{2}{L})$, the risk-sensitive cost admits the following bound:
@@
$\overline{R}_{K}^{(2)}(\theta)$ $\coloneqq\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K+1}}\left[\frac{1-\left[1-2\mu\alpha(1-\frac{L\alpha}{2})+\alpha\mu|1-\alpha L|s_{\scriptscriptstyle\mathrm{GD}}\right]^{K+1}}{2\mu\alpha(1-\frac{L\alpha}{2})-\alpha\mu|1-\alpha L|s_{\scriptscriptstyle\mathrm{GD}}}\left[f(x_{0})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\right]\right$.
@@
For NAG with $\alpha\in(0,1/L]$ and $\beta=\nu=\frac{1-\sqrt{\alpha\mu}}{1+\sqrt{\alpha\mu}}$, the bound (5.3) in theorem˜5.1 holds with constants
@@
$q=\frac{2}{s_{1}}\left(\frac{\alpha(1-\sqrt{\alpha\mu})}{4(1+\sqrt{\alpha\mu})^{2}}\right)+\frac{4L^{3}\alpha^{2}\nu^{2}}{2\mu s_{2}}\,,\qquad r=a+\alpha^{2}\left(\frac{c_{1}L}{2}+\tilde{P}_{11}\right)=s_{1}+\frac{L\alpha^{2}(2+s_{2})}{2s_{2}}+\frac{\alpha}{2},$
@@
where we used the definition of $\bar{c}(K+1,\theta)$ and $a_{j,K+1}$ given in theorem˜5.1. Since the sequence $\{a_{1,j}\}_{j\geq 1}$ is non-decreasing in $j$ admitting $\lim_{j\to\infty}a_{1,j}=\frac{1}{1-(p+q)}$; the sequence $\ell_{j}:=\log\left(\frac{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})+r~\theta a_{1,j}}{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})-r~\theta a_{1,j}}\right)$ is also non-decreasing and admits the limit
$\ell_{\infty}:=\lim_{j\to\infty}\ell_{j}=\log\left(\frac{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})+r~\theta 1/(1-(p+q))}{c_{1}+\frac{2}{L}\tilde{r}(\tilde{P})-r~\theta 1/(1-(p+q))}\right)=\log\left(\frac{1+\theta(\overline{H}_{\infty})^{2}}{1-\theta(\overline{H}_{\infty})^{2}}\right),$
using the definition of $\overline{H}_{\infty}$ from theorem˜5.1. Hence we conclude from (5.14) that
@@
Since $R(\theta)=\limsup_{K\to\infty}R_{K}(\theta)\leq\limsup_{K\to\infty}\overline{R}_{K}(\theta)$ and $R_{K}(\theta)=+\infty$ for $\sqrt{\theta}\,\overline{H}_{\infty}\geq 1$ for $\forall K$, we obtain (5.9) as desired. With similar computations, taking limit superior on both sides in (5.7) with respect to $K$ and using (5.8), we obtain $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq 4\sigma^{2}(\overline{H}_{\infty})^{2}.$ Finally, the fact that $\lim_{\theta\downarrow 0}\overline{R}(\theta)=4\sigma^{2}(\overline{H}_{\infty})^{2}$ follows simply by taking limit as $\theta\to 0$ in (5.9). ∎
@@
For momentum methods with constant step size and unbiased stochastic gradient estimates, ergodic averages converge to a neighborhood controlled by the variance proxy $\sigma^{2}$. For example, for NAG-beta-opt (using parameters as in table˜1) with i.i.d. noise $w_{k+1}\sim\mathcal{N}(0,\tfrac{\sigma^{2}}{d})$, [53, Cor. 4.8] gives $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\tfrac{\sqrt{\alpha}}{2\sqrt{\mu}}(1+\alpha L)$ for $\alpha\in(0,1/L]$, which vanishes as $\alpha\to 0$. In contrast, under biased noise the neighborhood is generally bounded away from zero, owing to a potentially non-vanishing worst-case drift. From (5.10), the asymptotic error is governed by the bound on $H_{\infty}$-norm. For NAG-beta-opt, under the ˜3.1, by corollary˜A.3 in the appendix, we have $\overline{H}_{\infty}=\tfrac{2\sqrt{5}}{\sqrt{\mu}}+\mathcal{O}(\sqrt{\alpha})$ as $\alpha\to 0$, and by (5.10) this yields $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\bigl(\tfrac{80}{\mu}+\mathcal{O}(\sqrt{\alpha})\bigr)$, and the right-hand side remains strictly positive as $\alpha\to 0$. Corollary 5.5 shows that momentum algorithms with smaller bound $\overline{H}_{\infty}$ exhibit tighter bounds on the asymptotic (suboptimality) bias, i.e., $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]$. Similar observations can be made for GD, HB or GMM. To our knowledge, this connection to the $H_{\infty}$-norm is entirely new for momentum methods.
For quadratic functions, we obtained large deviation results which obtains both the upper bound and the lower bound in (4.16) through an exact computation of the risk-sensitive index as $K\to\infty$. Beyond quadratics, we have only bounds on the risk-sensitive index. In the following result, we leverage this bound to obtain a large-deviation upper bound and high probability bounds for finite $K$.
@@
Under the premise of Theorem 5.1, consider the averaged iterates $\bar{x}_{K}=\frac{\sum_{j=0}^{K}x_{j}}{K+1}$. Then, for any $K\geq 1$ given and for any $\theta\in[0,~1/\overline{H}_{\infty}^{2})$ and $t\geq 0$, we have
@@
The first inequality in (5.15) is a consequence of the convexity of $f$. The second inequality is trivial for $\theta=0$, and for $(0,1/(\overline{H}_{\infty})^{2})$ it is simply a Chernoff-type bound. More specifically, for $t\geq 0$, and $\theta\in(0,1/(\overline{H}_{\infty})^{2})$,
@@
where we used $\overline{R}_{K}(\theta)\geq R_{K}(\theta)$. Then, (5.16) is obtained by taking supremum over $\theta$. Finally, the lower bound on $\bar{I}_{K}(t)$ given in (5.17) is a consequence of Lemma E.4 in the appendix.
@@
$\bar{I}(t){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=}\begin{cases}\displaystyle\frac{t}{2\sigma^{2}(\overline{H}_{\infty})^{2}}\sqrt{1-\frac{4\sigma^{2}(\overline{H}_{\infty})^{2}}{t}}-\log\left(\frac{1+\sqrt{1-\frac{4\sigma^{2}(\overline{H}_{\infty})^{2}}{t}}}{1-\sqrt{1-\frac{4\sigma^{2}(\overline{H}_{\infty})^{2}}{t}}}\right),&\text{if}\qquad t\geq 4\sigma^{2}(\overline{H}_{\infty})^{2};\\ 0,&\text{if}\qquad t<4\sigma^{2}(\overline{H}_{\infty})^{2}.\end{cases}$ (5.20)
@@
By Corollary 5.5, $\limsup_{K\to\infty}\,\mathbb{E}\!\left[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\right]=\overline{m}:=4\sigma^{2}(\overline{H}_{\infty})^{2}.$ Since this value represents the asymptotic expected suboptimality, the events with suboptimality below $\overline{m}$ are not exponentially unlikely; hence, $\bar{I}(t)=0$ whenever $t<\overline{m}$, in accordance with (5.20). In contrast, for $t>\overline{m}$, (5.20) shows that $\bar{I}(t)$ is strictly decreasing in $\overline{H}_{\infty}$ over the region where $\bar{I}(t)>0$, i.e., when $0<\overline{H}_{\infty}<\tfrac{\sqrt{t}}{2\sigma}$. Therefore, smaller values of $\overline{H}_{\infty}$ correspond to strictly larger decay rates $\bar{I}(t)$ at the threshold $t$, provided that $t$ exceeds the critical level $\overline{m}=4\sigma^{2}(\overline{H}_{\infty})^{2}$. This shows that a smaller bound on $H_{\infty}$-norm—reflecting greater robustness to worst-case deterministic noise—leads to sharper concentration properties and faster decay of the tail probabilities associated with $f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}$. This is broadly consistent with the numerical behavior observed in LABEL:fig:rate-fn-nag-beta-opt-vary-stepsize for quadratics: reducing the step size increases the rate function $I(t)$ while simultaneously decreasing the bound on $H_{\infty}$-norm for NAG-beta-opt, implied by (A.8) in the appendix.